OpenStudy (anonymous):

y varies jointly with p and q. When p = 10 and q = 2, y = 60. Find y when p = 5 and q = 5.

OpenStudy (cwrw238):

y = kpq where k is a constant plug in y=60.p=10 and q=2 to find k then you can use this for last part of your question.

OpenStudy (anonymous):

i dont get it

OpenStudy (cwrw238):

y = kpq = k*p*q 60 = k *10*2 do u get that?